OpenStudy (anonymous):

Integral of x^2 dx/x^2 + 9 Use partial fraction decomposition, and long division if needed.

OpenStudy (solomonzelman):

I will post an example

OpenStudy (solomonzelman):

For example: \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x^2}{x^2+4}~dx}\) You will perform this trig-sub: \(\large\color{black}{\displaystyle x=\sqrt{4}\tan\theta=2\tan\theta}\) \(\large\color{black}{\displaystyle dx/d\theta =2\sec^2\theta\quad\Longrightarrow\quad dx=2\sec^2\theta{~}d\theta }\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{\left(2\tan\theta\right)^2}{\left(2\tan\theta\right)^2+4}~\left(2\sec^2\theta{~}d\theta\right)}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{4\tan^2\theta}{4\tan^2\theta+4}~\left(2\sec^2\theta{~}d\theta\right)}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{8\tan^2\theta\sec^2\theta}{4\tan^2\theta+4}~{~}d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{8\tan^2\theta\sec^2\theta}{4(\tan^2\theta+1)}~{~}d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{8\tan^2\theta\sec^2\theta}{4(\sec^2\theta)}~{~}d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}2\tan^2\theta{~}d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}2\sec^2\theta+2{~}d\theta}\) \(\large\color{black}{\displaystyle 2\tan\theta+2\theta+C}\) Now we will rearrange \(\theta\) in terms of x: \(\large\color{black}{\displaystyle x=2\tan\theta}\) \(\large\color{black}{\displaystyle \frac{x}{2}=\tan\theta}\) \(\large\color{black}{\displaystyle \theta=\arctan\left(\frac{x}{2}\right)}\) Now do the substitution: \(\large\color{black}{\displaystyle 2\tan\left(\arctan\left(\frac{x}{2}\right)\right)+2\arctan\left(\frac{x}{2}\right)+C}\) \(\large\color{black}{\displaystyle 2\cdot\frac{x}{2}+2\arctan\left(\frac{x}{2}\right)+C}\) \(\large\color{black}{\displaystyle x+2\arctan\left(\frac{x}{2}\right)+C}\)

OpenStudy (solomonzelman):

I just used a 4 instead of a 9, but I hope that this helps.

OpenStudy (anonymous):

Trig substitution, yeah... I'm just not sure if I'm supposed to solve the problem with just the long division/partial fractions. I guess not, if you're not seeing a way to do it that way.

OpenStudy (solomonzelman):

I don't think you can do partial fractions on this because the bottom can not be factored. Trig sub will be needed

OpenStudy (solomonzelman):

\(\rm x^2+4\) is not factorable (well, not into reals)

OpenStudy (anonymous):

Right. Thanks!

OpenStudy (solomonzelman):

You Welcome !

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

this is wrong

OpenStudy (solomonzelman):

Are you sure that this is wrong? I really doubt that there is any error in my example, although anyone can make a mistake. If you think that I really made a mistake, then please pint out where and how.

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

nooooooooooooooo

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

you're fine

OpenStudy (solomonzelman):

Then what is wrong? To do what I did is a violation of the site's policy?

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

wat o.O

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

my love for you is wrong ;-; k bye

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{a \rightarrow ~\infty}\left( x{\rm D} \right)^a}\)

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

awh :') ofc our love is beyond infinity <3